OpenStudy (anonymous):

2x + 5 = 10 What is the first operation used in solving this equation? A) addition B) division C) multiplication D) subtraction

OpenStudy (anonymous):

D

OpenStudy (anonymous):

You have to get rid of the 5. So you subtract it from the 10.

OpenStudy (anonymous):

i thought it was that..thanks..can u help me on somemore

OpenStudy (anonymous):

just use pemdas backwards (:

OpenStudy (anonymous):

Yea . WHat are the other ones?

OpenStudy (anonymous):

subtraction coz u will be transposing 5 to RHS so it will become -5 so 10-5

OpenStudy (anonymous):

Add -4 + 2/3 . A) -5 2/3 B) -4 2/3 C) -3 2/3 D) -3 1/3

OpenStudy (anonymous):

D

OpenStudy (anonymous):

r u sure..an di have more also..thanks soo much..yall r a big help

OpenStudy (anonymous):

Yup , Its D

OpenStudy (anonymous):

kk cool

OpenStudy (anonymous):

Anything else ?

OpenStudy (anonymous):

yes im sure (:

OpenStudy (anonymous):

2(x + 7) + 3x = 12 What property is used in the first step in solving this equation for x? A) Reflexive Property B) Transitive Property C) Commutative Property D) Distributive Property

OpenStudy (anonymous):

D again lol

OpenStudy (anonymous):

you have to distribute the 2 to the x and the 7

OpenStudy (anonymous):

kk thanks..i got it.. :] 18) What is the value of x, when -25(x - 4) = -55(x - 10)? A) x = 5 B) x = -5 C) x = 8 1 8 D) x = 15

OpenStudy (anonymous):

D again! x= 15 lol

OpenStudy (anonymous):

-25(x - 4) = -55(x - 10) distribute..... -25x+100 = -55x+550 -100 -100 subtract from both sides.... -25x=-55x+450 +55x +55x add to both sides.... 30x=450 --------- 30 divide by 30 x=450

OpenStudy (anonymous):

x=15* oops

OpenStudy (anonymous):

lol kk thanks OK Rolls Rejected Rolls Rolls to Repair 75 12 8 % of Total % of Total % of Total A B C A carpet inspector was inspecting short rolls of carpet. The summary of her inspection is shown in the table. She needs to complete the table, which of the choices shown is a correct percentage to fill in the table? A) A = 78.9% B) B = 16% C) B = 8.42% D) C = 9.2%

OpenStudy (anonymous):

A= 75 is 78.94 percent of 95 B= 12 is 12.63 percent of 95 C= 8 is 8.42 percent of 95 so the correct answer is A

OpenStudy (anonymous):

kk thanks..i was close..lol

OpenStudy (anonymous):

5x + 7 < 3x + 3 (given) 2x + 7 < 3 (subtraction) 2x < -4 (subtraction) x < -2 ( ? ) Fill in the missing reason for the argument shown. A) addition B) division C) given D) subtraction

OpenStudy (anonymous):

B division. They divided both sides by 2 to get the x by itself. -4/2=-2 (:

OpenStudy (anonymous):

thanks so much fro explaining it also lol Johnson High School has a total enrollment of 1829 students. Season tickets are sold for the basketball season and 604 students buy the tickets. About what percent of the student body purchased season tickets? A) 3% B) 25% C) 33% D) 40%

OpenStudy (anonymous):

@indigochild ..THIS IS MY LAST ONE PLZZ

OpenStudy (anonymous):

lol no problem! 604 is around 33 percent of 1829, so it would be C (:

OpenStudy (anonymous):

let me know how you did!

OpenStudy (anonymous):

kk cool..r u sure your answers are correct..i really need this

OpenStudy (anonymous):

yes im pretty sure (:

OpenStudy (anonymous):

yep thanks..all of you answers we correct..some of the ones i answerd werent so now iget 2 retry 3 times..i have to make over a 70 and i made a 69

OpenStudy (anonymous):

try again, ill help you (:

OpenStudy (anonymous):

kk thanks.. The Bergstedts are going on a family vacation. Mrs. Bergstedt needs to make sure she has enough movies for the car ride, which will be 8.5 hours. She has a total of 6.8 movies (the last videotape ran out before it was finished). She divides 8.5 by 6.8 and gets 1 with a remainder of 25. What does that mean? A) That Mrs. Bergstedt needs an additional 1.25 movies to cover the trip. B) That Mrs. Bergstedt needs 1.25 movies for her to have enough to cover the entire trip. C) There is no real world meaning as you cannot divide hours by movies and have it mean anything. D) That in order for the kids to watch movies the entire trip each one needs to be 1.25 hours long.